OpenStudy (anonymous):

state any excluded values m-3/3m-9

OpenStudy (anonymous):

I need help!!!:(

jhonyy9 (jhonyy9):

how you know when is a fraction undefined ?

OpenStudy (anonymous):

uhhi dont know haha. Im really bad at this :/ sorry

jhonyy9 (jhonyy9):

so when the denominator is equal zero so than you need getting the value of m for what the denominator will be equal zero

jhonyy9 (jhonyy9):

for you make 3m-9=0 3m=9 m=9/3 m=3 so for m=3 this fraction is undefined

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

thanks so much! :)

jhonyy9 (jhonyy9):

yw was my pleasure good luck bye

OpenStudy (anonymous):

y u block? thts messed up?